OpenStudy (anonymous):

How would you calculate the number of moles and atoms in a sample of gold that had a mass of 35.12 grams?

OpenStudy (anonymous):

1 mole=6022*10 to the 23 power of molecules (: So, just flip it. :D

OpenStudy (anonymous):

@jaderoman Have you got the answer?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Ok. First, we need to find the moles of gold

OpenStudy (anonymous):

Next, one mole of gold contains 6.022x10^23 Au atoms. Therefore, if we multiply moles of Au by 6.022x10^23, we will get the number of Au atoms.

OpenStudy (anonymous):

\(\sf 35.12g\times \large {\frac{1~mole~Au}{196.97~g}} \times \large {\frac{6.022\times10^{23}~Au~atoms}{1~mole~Au}} = ...Au~atoms\)

OpenStudy (anonymous):

If you don't understand, I will try to explain

OpenStudy (anonymous):

I dont understand.. i'm just really confused with this whole topic

OpenStudy (anonymous):

@tinybookworm could you please explain

OpenStudy (anonymous):

Sorry for letting you wait

OpenStudy (anonymous):

First, you may know that \(\sf mass=moles\times molar~mass\), right?

OpenStudy (anonymous):

@jaderoman Are you there?

OpenStudy (anonymous):

yes sorry I haven't been on

OpenStudy (4n1m0s1ty):

I guess i'll continue on from tinybookworm. What are you confused on?

OpenStudy (anonymous):

I just don't know how to do it to be honest. I've been really behind on my schoolwork so I haven't been doing too well with it.

OpenStudy (4n1m0s1ty):

Ok, so lets start from the beginning. The problem asks for the number and moles/atoms in a sample of gold. The only information it gives you is the mass of the sample of gold.

OpenStudy (4n1m0s1ty):

So lets define these terms, do you know what mass, moles, and atoms are?

OpenStudy (anonymous):

Yes

OpenStudy (4n1m0s1ty):

Ok good

OpenStudy (4n1m0s1ty):

Do you what molar mass is? (its sometimes also called atomic weight/atomic mass)

OpenStudy (anonymous):

Sorry I hope you're still there, my wifi went out. Anyways, isn't it the amount of something in a substance?

OpenStudy (4n1m0s1ty):

It describes the amount of mass per mole of a substance. If you look on the periodic table, its ypically the bottom number on the element.

OpenStudy (anonymous):

Ohh okay

OpenStudy (anonymous):

So I need to find the molar mass for Au?

OpenStudy (4n1m0s1ty):

Yes

OpenStudy (anonymous):

Its 196.966543?

OpenStudy (4n1m0s1ty):

Its given on the periodic table, you just need to look for it.

OpenStudy (4n1m0s1ty):

Yes

OpenStudy (anonymous):

okay so what do i do with that

OpenStudy (4n1m0s1ty):

The entire meaning of that number is telling you that 1 mole of Au is equivalent to having 196.97 g of Au

OpenStudy (anonymous):

okay so does that mean i would divide 35.12 by that? to find the amount of moles it is?

OpenStudy (4n1m0s1ty):

So I want to use an analogy so you can understand this better. Say that I have a group of cars, and each of the cars of 2 people in them. If I have 8 cars, how many total people would I have? I would have 16 people in total right? I just took the number of cars and multiplied by the number of people per car. Now, lets take this in reverse, if I had 20 people, how many cars would I have? Its 10 cars. I just took the amount of people and divided by the number of people per car, in order to find the total number of cars. Anyways, this car analogy is pretty much what you are doing for this chemistry (stochiometry) problem. So just think about that for a bit, while I help someone else real quick. I'll be back in a minute or two.

OpenStudy (anonymous):

okay.. that just confused me

OpenStudy (anonymous):

what are the cars and people supposed to represent?

OpenStudy (4n1m0s1ty):

The car and people are like mass and moles/atoms. While the number of people per car is like the molecular mass.

OpenStudy (anonymous):

oh okay that makes sense

OpenStudy (4n1m0s1ty):

So in order to convert between mass and moles/atoms. We need to use molecular mass as kind of an intermediary value, like I did in the car example.

OpenStudy (4n1m0s1ty):

So since I have the mass of Au, in order to get moles, I need to do this: \[\frac{35.12\,g\,Au}{}*\frac{1\,mole\,Au}{196.97\,g\,Au} = ?\,mole\,Au\] So I need to divide by the molecular weight in order to get the number of moles. I setup the problem in the fractional form above, so its a bit more explicit about what I'm really doing. I'm essentially using the fraction, so I can figure out whether to divide or multiply by the molecular weight. I'm dividing, so I can eliminate the mass unit (grams) and I'm left with moles as my answer.

OpenStudy (4n1m0s1ty):

To find the number of atoms you just need to know that 1 mole = 6.022 x 10^23 atoms. Then like what we did with the molecular weight above, use that as an intermediary value to convert from moles to atoms. So like this: \[\frac{?\,moles\,Au}{}*\frac{6.022*10^{23}\,atoms}{1\,mole} = ?\,atoms\,Au\] So the fraction tells me to multiply by the intermediary value to get atoms.

OpenStudy (anonymous):

so for the first one i would divide 35.12 by 196.97?

OpenStudy (4n1m0s1ty):

Yep

OpenStudy (4n1m0s1ty):

and the second step you multiply.

OpenStudy (4n1m0s1ty):

But does it make sense to you how I know when to multiply/divide?

OpenStudy (anonymous):

so i guess i would say about 0.18 moles? and i think i get it

OpenStudy (anonymous):

and i plug that into the second equation?

OpenStudy (4n1m0s1ty):

Yep, I got 0.18 moles, and you plug that value into the first ? mark I made in the second equation.

OpenStudy (anonymous):

I'm kind of confused on what I'm multiplying though

OpenStudy (4n1m0s1ty):

On the second equation? We are multiplying by the number of atoms in 1 mole of a substance. Thats what 1 mole is defined as, 6.022 * 10^23 units

OpenStudy (anonymous):

so 0.18 x 6.022 x 10^23?

OpenStudy (4n1m0s1ty):

Yes

OpenStudy (anonymous):

is it supposed to be a really big number?

OpenStudy (anonymous):

or did i do it wrong

OpenStudy (4n1m0s1ty):

yep, it should be. A mole is 6.022*10^23, which is about the same size as the volume of water in the pacific ocean. By multiplying by 0.18 you are essentially only taking 18% of that value. So think about 18% of the pacific ocean, thats still really big.

OpenStudy (4n1m0s1ty):

1.08*10^23 atoms Thats what I got.

OpenStudy (anonymous):

Okay do you think you could help me with the other problems too?

OpenStudy (4n1m0s1ty):

Sure, I would start a new thread though. I'm currently helping someone else at the moment, but I'll try to help you out as soon as I'm done. If I don't get to you in time, you might be able to get help from someone else.

OpenStudy (anonymous):

it's basically the same thing except i have to do it with table sugar, which has 3 different elements in it so I'm not really sure what to do there

OpenStudy (4n1m0s1ty):

Well instead of just one atom you have a multiple of atoms in a molecule. The only thing different is that you are dealing with molecular mass rather than atomic mass.

OpenStudy (anonymous):

Okay I'll just have to start a new thread because I'm not quite sure how to set it up

OpenStudy (anonymous):

@4n1m0s1ty Thank you for your explanation